OpenStudy (anonymous):

Please help me with my math preferably with step by step. (4x+4)^1/2-3x=0 I am not good at math at all :( so please dumb it down as much as possible :3

OpenStudy (anonymous):

Nah it seems like \[ (4x+4)^{\frac{1}{2}}-3x=0\]

OpenStudy (anonymous):

which is \[ \sqrt{4x+4}-3x=0\]

OpenStudy (anonymous):

ummm are we solving for x or are we simplifying?

OpenStudy (ranga):

Solving for x. Isolate the square root on the left. Square both sides to get rid of the square root. Then bring everything to the left side. You will have a quadratic equation. Solve for x. You will get two values. Put them back into the equation to make sure they are correct.

OpenStudy (anonymous):

@BlackLabel on my work it says solve. the possible answers are A. -4/9, B. 1, C. -4/9,1 or D. no solution

OpenStudy (anonymous):

Well Follow what @ranga said. The first step would be to bring the 3x to the other side of the equation

OpenStudy (anonymous):

And then we would square both sides

OpenStudy (ranga):

Are you sure you typed the question correctly? None of the answers seem to satisfy the equation. Perhaps you can attach a picture of the question. Is the 4x+4 supposed to be 5x + 4?